A large cooperation has quality control over its fertilizers. The fertilizes are composed of nitrogen. The fertilizer requires 3 mg of nitrogen. The distribution of the percentage of nitrogen is unknown with a mean of 2.5 mg and a standard deviation of 0.1. A specialist randomly checked 100 fertilizer samples. What is the probability that the mean of the sample of 100 fertilizers less than 2 mg?
Answer #1

Given,

\mu = 2.5 , \sigma = 0.1

Using central limit theorem,

P(\bar{x} < x) = P(Z < (x - \mu ) / (\sigma / sqrt(n) ) )

So,

P(\bar{x} < 2) = P(Z < (2-2.5) / (0.1 / sqrt(100 ) ))

= P(Z < -50)

= 0 .0000